OpenStudy (anonymous):

find the area in oval limacon r=5+cos(theta)

OpenStudy (anonymous):

it's a circle with radius 5.

OpenStudy (anonymous):

need the area

OpenStudy (anonymous):

actually my bad, doesn't look like a circle though.

OpenStudy (anonymous):

okk

OpenStudy (anonymous):

anyway, area = integral (1/2) (5 + cos(theta))^2 dtheta, from 0 to 2pi

OpenStudy (anonymous):

try \[\int\limits_{0}^{2\pi} 0.5(5+\cos \theta)^2d\]

OpenStudy (anonymous):

do the math and area = (51/2)pi

OpenStudy (anonymous):

which is what she said

OpenStudy (anonymous):

yess thank you